OpenStudy (anonymous):

T=2(3.14)√1.6/32 . Solve. Can someone explain how I can solve this?

OpenStudy (phi):

This is an order of operations question: pema parentheses, exponents (and square roots), multiplication (and division), addition (and subtraction) However, for the specific answer we need to know what is under the square root: is it \[ T= 2(3.14)\sqrt{\frac{1.6}{32}} \]? if so, use a calculator and do it in this order: 1.6/32 = 0.05 take the square root of 0.05 = 0.2236 multiply by 3.14 = 0.7021 multiply by 2= 1.404